OpenStudy (anonymous):

If someone can just like... explain this to me it would be a huge help. Im so lost. 1/2+2/3

OpenStudy (anonymous):

first step is to make the denominators equal. in this case, you would multiply the denominators by each other's fraction(so that the nominator is evened out). \[{1 \times 3 \over 2 \times 3} + {2 \times 2 \over 3 \times 2}\]

OpenStudy (anonymous):

\[{3 \over 6} + {4 \over 6}\]

OpenStudy (anonymous):

Okay.. then what?

OpenStudy (anonymous):

Add the two fractions together. what do you get?

OpenStudy (anonymous):

7/6?

OpenStudy (anonymous):

or 7/12...

OpenStudy (anonymous):

your first answer was correct. Denominators don't change during addition.

OpenStudy (unklerhaukus):